a recipe calls for 2 1/3 cups broth. A cook is making 3/4 of the recipe. what IS THE AMOUNT OF BROTH NEEDED?
Answer by hovuquocan1997(83)

Now you have two convert the mixed number fraction 2 1/3 into a normal fraction first, you will have
2 1/3 = 7/3 (you can also use calculator)
so a recipe calls for 7/3 cups broth, but a cook is only making 3/4 of the recipe. So you will have
7/3 * 3/4 = 7/4
TA-DAH
You only need 7/4 (or you can write 1 3/4) cup of broth :D